OpenStudy (anonymous):

Y (-1)*(-1) = 1 Due to this we have imaginary no.s................

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

it's not a question It's a paradox

OpenStudy (turingtest):

paradox?? how so?

OpenStudy (anonymous):

Do u know iota y we call it imaginary coz of this Y -1 * -1 = -1

OpenStudy (turingtest):

what's the point? and what is the 'Y' ? and why did you write -1*-1=-1 ? ^^^not true

OpenStudy (anonymous):

First tl me do u know complex no.s rather imaginary one

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

Tl me Y iota is imaginary to our mathematics Y she can't imagine iota

OpenStudy (anonymous):

The reason for the fallacy is that sqrt(-1) is not an ordinary (real) square root, hence the rule for computing the quotient of radicals does not apply to it. and this, kids, is what happens when you don't know how to read

OpenStudy (anonymous):

im sure thats what he means

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I mean if we develop axioms for mathematics then y we didn't choose -1 * -1 = -1 Coz if this is so then sqrt(-1) = -1

OpenStudy (anonymous):

gee, this guy is a genius, how could we miss that

OpenStudy (turingtest):

of course that's not an axiom, its a consequence of the real axioms. If we adopted that one math would be inconsistent. I don't think you have a good feel for axiomatic formulation.

OpenStudy (anonymous):

if (-1)*(-1)=-1 then\[-8=8(-1)(-1)=-8(-1)=???\]

OpenStudy (anonymous):

It must be having reason Is anyone doing Bachelors or masters or phd in maths

OpenStudy (anonymous):

im doing postdoc and have fields medal and READ what turing wrote

OpenStudy (anonymous):

It's not easy to say this we adopted

OpenStudy (anonymous):

Mathematicians fear from -ve no.s till 20th century

OpenStudy (anonymous):

1/0=0 i solved the world

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

1/0=1/1*0/0=1*0=0

OpenStudy (anonymous):

it's wrong u don't know 0 dear

OpenStudy (anonymous):

u dont know -1

OpenStudy (anonymous):

loooool

OpenStudy (anonymous):

that's y i was asking

OpenStudy (anonymous):

K leave this tell me How & what & y is pie

OpenStudy (anonymous):

does this have to do with -1 too?

OpenStudy (anonymous):

No I hope to find Y things are so U are doing postdoc na From where

OpenStudy (anonymous):

pre-algebra, could you please sort out whether pi is infinite or not, ever recurring, and all other unknowns associated with it, and figure out more than 5 trillion digits before the day's out? Cheers bud.